determine whether the graphs of the lines are parallel and show your work.
x+9=y and y-x=-5

Answer by grishma.kshatriya(27)

Answer :
First we write both the equation in Y-intercept form i.e y = mx + b .
x + 9 = y
y = x + 9......(1)

y - x = -5
Adding x on both the sides we get ,
y = x - 5......(2)
Comparing (1) and (2) to y = mx + b and find the slope m
We get m1 = 1 and m2 = 1
We know that if the slopes of two lines are same then they are said as parallel lines.
So given two lines are parallel .
Thank you .